OpenStudy (anonymous):

Solve the equation. Express radicals in simplest form. 2m^2+12m+1=0

OpenStudy (anonymous):

Answer=( -12+136^1/2)/4=-3 +(34^1/2)/2 and also -3 - (34^1/2)/2

OpenStudy (anonymous):

\[m=\frac{-12 \pm \sqrt{12^2-4(2)(1)}}{2(2)}\] \[\implies m=\frac{-12\pm 2\sqrt{34}}{4}\]

OpenStudy (gw2011):

Use the quadratic formula: x=[-b+/-square root (b^2-4ac)]/2a a=2 b=12 c=1 x=m m=[-12+/-square root (12^2-4*2*1)]/(2)(2) m=[12+/-square root (144-8)]/4 m=[-12+/-square root (136)]/4=[12+/-square root 4*34]/4 m=[-12+/-(2*square root 34)]/4 m=[-6+/-(square root 34)]/2 m=(-6+square root 34)/2 and m=(-6-square root 34)/2